OpenStudy (anonymous):

Jen is 13 years younger than Andre. the sum of their ages in years is 137. What is andre's age in years

OpenStudy (anonymous):

J + 13 = A J + A = 137 Substitute: J + (J + 13) = 137 2J + 13 = 137 2J = 124 J = 62 Plug this back in to the original: J + 13 = A 62 + 13 = A A = 75 Jen is 62, and Andre is 75 I found this on Yahoo! But Im pretty sure its right.